You have space in your garden for 40 small flowering bulbs. Crocus bulbs cost $0.55 each and daffodil bulbs cost $0.65 each. Your budget allows you to spend $23.20 on bulbs. How many crocus bulbs and how many daffodil bulbs can you buy?
Answer by Cromlix(4381)

Crocus = x
Daffodil = y
x + y = 40........1
55x + 65y = 23.20.....2
Multiply 1 by 65
65x + 65y = 2600.....3
55x + 65y = 2320.....2
Subtract (2) from (3)
10x = 280
x = 28
Substitute x = 28 in
Eq (1)
x + y = 40
28 + y = 40
y = 40 - 28
y = 12
28 Crocus and 12 Daffodil
Hope this helps.
:-)